# ReNiL: Relative Neural Inertial Locator with Any-Scale Bayesian Inference

### v2

@@ IV-A Orientation Alignment
IV-B ASLE Structure
We propose an any-scale Laplace estimator (ASLE), which consists of five key components: 1) the patching module, 2) the embedding module, 3) the feature extractor, 4) the contextual builder, and 5) the PCO module. Each module within the ASLE can be replaced with other general modules that perform the same function. In this paper, we have designed a specific verification model for the system. The structure of the ASLE is shown in Fig. 3.
@@ IV-B ASLE Structure
Embedding module
The embedding module is used to map the original 6 features of the input $X\in\mathcal{R}^{B\times P\times 6\times L}$ to a higher-dimensional space and compress the patch length to greatly reduce computation overhead. Specifically, a TCN compressor is applied to embed $X$ from 6 dimensions into a higher dimension and L into a shorter length:
@@
where $e$ is the embedding dimension and $L^{\prime}$ is the compressed length. Finally, a ReLU layer and a group normalization (GN) layer are applied for activation and normalization, which helps to speed up the network convergence.
@@
The 1D version of ResNet18 [35] has demonstrated strong sequence feature extraction capabilities in previous work. We use its improved residual block (replace BN with GN) as a feature extractor to capture motion pattern features in each patch in parallel. Specifically, for $X^{e}$, multiple 1D residual blocks are applied sequentially to extract patch features. Each residual block consists of two convolutional layers and two GN layers. The number of channels in each residual block is double that of the previous block. The module contains a total of $N_{1}$ residual blocks. For each block, the convolutional layer has a stride of 1 and a kernel size of 3. The output of each 1D residual block serves as the input for the next 1D residual block. The outputs of the final residual block are concatenated along the channel axis. The output of this module is:
@@
2D Res block is highly effective for local context modeling due to its ability to expand the receptive field as the number of layers increases. It also offers fast computation, making it an excellent choice for modeling movement context between patches. Specifically, to allow the convolution kernel to compute patches and features in the last two dimensions, we rotate the second and third axes of the output tensor $X$ from the 1D feature extractor. This yields a new tensor
$X\in\mathcal{R}^{B\times C\times P\times W}$ (16)
Then, $N_{2}$ 2D residual blocks are applied to build the temporal local context of the features, resulting in a robust features map. The number of channels in each residual block is the same as in the previous block. Each 2D residual block consists of two convolutional layers and two GN layers, with a kernel size of $3\times 3$ and a stride of 1 for each convolutional layer. Finally, we obtain robust features map:
$X\in\mathcal{R}^{B\times C\times P\times W}$ (17)
This operation aims to mix the contextual potential relationships between patches and improve the robustness of the final features map.
PCO module
The scale pooling, concatenation, and output module aims to capture both the common features inherent in each patch, the contextual information between patches, scale-aggregate them into a unified dimension, and regress. Specifically, multi-kernel adaptive average pooling operations are applied to the second and fourth axes of the output from the feature extractor to obtain the inherent features $F_{1}\in R^{B\times C_{1}}$ between different patches. Multi-kernel adaptive average pooling operations are applied to the last two axes of the contextual builder output to obtain the inherent features $F_{2}\in R^{B\times C_{2}}$. These unified features are concatenated together:
$X=F_{1}\oplus F_{2},X\in R^{B\times(C_{1}+C_{2})}$ (18)
And regress $\Delta\hat{p},\hat{b}$ through the output layer $O(\cdot)$, which contains fully connected, dropout, and ReLU layers.
@@
IV-C Joint Training with Uncertainty Estimation
@@
Specifically, for the original input $X\in R^{B\times P\times 6\times L}$ of the network, we record its features map $X_{f}\in R^{B\times C\times P\times W}$, which is the output after the contextual builder. We then perform a series of data augmentations on the original input $X$, including:
@@
Subsequently, we obtain the features map $\hat{X}_{f}\in R^{B\times C\times P\times W}$ of the contextual builder output. Finally, we calculate the MSE loss $L_{fm}$ between $\hat{X}_{f}$ and $X_{f}$ to construct robust context modeling features:
@@
IV-D Bayesian Process
The output of ASLE is the relative displacement and scale parameters between two IPDPs. However, our localization inference is a continuous process. Next, we will use Bayesian inference to connect all IPDPs in series to form a complete Bayesian inference chain. In the case of IPDPs involved with other external localization sources, we use a Bayesian filter [36] for loose coupling.
@@ IV-D Bayesian Process
$\mathbf{p}\left({p}_{t_{n}}\mid u_{t_{0}:t_{n}}\right)$ $=\int\mathbf{p}\left({p}_{t_{n}}\mid{p}_{t_{n-1}},u_{t_{n}}\right)$ (27)
$\mathbf{p}\left({p}_{t_{n-1}}\mid u_{t_{0}:t_{n-1}}\right)d{p}_{t_{n-1}}$
@@
$\mathbf{p}\left(\widetilde{p_{t_{n}}}\mid{z}_{t_{0}:t_{n-1}},u_{t_{0}:t_{n}}\right)=$ (29)
$\int\mathbf{p}\left(p_{t_{n}}\mid p_{t_{n-1}},{u}_{t_{n}}\right)\,\mathbf{p}\left(p_{t_{n-1}}\mid{z}_{t_{0}:t_{n-1}},{u}_{t_{0}:t_{n-1}}\right)\,dp_{t_{n-1}}$
@@
V Experimental Evaluations
V-A Experiment Setup
@@
For displacement-regression experiments we employ ASLE-ns, a family of our ASLE that accepts fixed n-second input windows. Because the competing baselines PDR, RoNIN-ResNet18 (RoRes18), TLIO, CTIN, and iMoT can each operate only at a single preset window size, we instantiate four representative variants (n = 1, 5, 10, 20) to span short, medium, and longer motion segments; these values were chosen simply to give broad temporal coverage rather than for any dataset-specific reason. We evaluate all methods head-to-head and then carry forward the two strongest baselines for deeper analysis. For device orientation estimation, our motion-aware orientation filter is benchmarked against the classical Mahony [37] and Madgwick [38] algorithms.
@@
The WUDataset collection was carried out using a Google Pixel 6A smartphone (IMU sampling rate: 200Hz), a Vicon optical system (sampling rate: 100Hz), and a LiDAR-SLAM system (sampling rate: 200Hz), all synchronized through a custom-developed Android application. The WUDataset is organized as follows:
@@
TABLE I: Evaluation Metrics and Definitions
@@
MAE $\mathrm{MAE}=\frac{1}{N}\sum_{i=1}^{N}|\hat{p}_{i}-p_{i}|$
ADE $\mathrm{ADE}=\frac{1}{N}\sum_{i=1}^{N}\frac{\|\Delta\hat{p}_{i}-\Delta p_{i}\|_{2}}{\Delta t_{i}}$
HE $\mathrm{HE}=\frac{1}{N}\sum_{i=1}^{N}|\hat{\theta}_{i}-\theta_{i}|$
QAE $\mathrm{QAE}=\frac{1}{N}\sum_{i=1}^{N}2\arccos(|\langle\hat{\mathbf{q}}_{i},\mathbf{q}_{i}\rangle|)$
CS $\mathrm{CS}=\frac{1}{N}\sum_{i=1}^{N}\frac{\langle\hat{\mathbf{q}}_{i},\mathbf{q}_{i}\rangle}{\|\hat{\mathbf{q}}_{i}\|\|\mathbf{q}_{i}\|}$
V-B Implementation Details
@@
TABLE II: Displacement accuracy performance on RoNIN-ds and WUDataset. Top 3 results in each row are marked with red, blue, green.
Dataset Subjects Metric PDR RoRes18 TLIO CTIN iMoT ASLE-1s ASLE-5s ASLE-10s ASLE-20s
@@
ASLE details
We implemented ASLE using the PyTorch 2.5.1 [39] framework on a server equipped with an Intel Xeon Silver 4210R CPU and a GeForce GTX 3090 GPU. The model utilizes 200Hz IMU input, with a patch length of 100. The embedding channel, kernel size, and stride of the two-layer TCN compressor are set to 32, 3, 3, and the GN size is set to 4. The feature extractor consists of four 1D residual layers with output channels of 64, 128, 256, and 512, respectively. The context builder consists of two 2D residual layers with output channels of 512 for each layer. The PCO scale pooling layer is conducted as three 2D adaptive average pooling with output size $1\times 1,2\times 1,3\times 1$ and a 2D adaptive max pooling with output size $3\times 1$ for contextual builder output, and a 2D adaptive average pooling with output size $1\times 1$ for feature extractor output. Then PCO passes the features through the fully connected layer with hidden size 1024, ReLU layer, and the output is a vector of length 4. Adam [40] is used as the optimizer, with the learning rate scheduler set to ReduceLROnPlateau [41], using a factor of 0.1 and patience of 10. The initial learning rate for the network parameters is set to $1\times 10^{-4}$, with a minimum learning rate of $1\times 10^{-12}$. The mini-batch size is 128, and the model is trained for a total of 200 epochs.
Bayesian process
$\tau_{t_{n}}\sim\operatorname{Exp}(\frac{1}{2\boldsymbol{\Sigma}_{b}^{2}})$ (31)
where $\boldsymbol{\Sigma}_{b}=\operatorname{diag}\left(\Delta t_{n}b_{v_{x}},\Delta t_{n}b_{v_{y}}\right)$. So the Laplace distribution noise can be conditionally Gaussian representation into a Gaussian distribution noise:
${w}_{t_{n}}\mid\tau_{t_{n}}\sim\mathcal{N}\left({0},\tau_{t_{n}}\boldsymbol{\Sigma}_{w}\right)$ (32)
where $\boldsymbol{\Sigma}_{w}=\operatorname{diag}\left(2\left(\Delta t_{n}b_{v_{x}}\right)^{2},2\left(\Delta t_{n}b_{v_{y}}\right)^{2}\right)$. Then, we use $\tau_{t_{n}}\boldsymbol{\Sigma}_{w}$ to construct a covariance matrix that conforms to the Gaussian distribution to complete the Bayesian inference chain.
$\delta_{t_{n}}=\left(p_{t_{n-1}}-p_{t_{n}}\right)^{\top}\boldsymbol{\Sigma}_{w}^{-1}\left(p_{t_{n-1}}-p_{t_{n}}\right)$ (33)
V-C Experimental Results of ASLE Network
Displacement accuracy performance
The purpose of this experiment is to evaluate the displacement accuracy of the ASLE network compared to several baseline methods, tested on both seen and unseen subjects in the RoNIN-ds and WUDataset datasets. As shown in Table II, when limited to a fixed time scale of 1 second, the displacement accuracy of all methods remains relatively similar across both datasets. When the time scale restriction is removed, the distinctive advantage of ASLE becomes evident. Its inherent design to process variable-scale IMU sequences allows it to outperform other methods that rely on fixed-length sliding windows. This flexibility enables ASLE to effectively capture temporal dependencies, which translates into substantially improved displacement accuracy, particularly pronounced on the more complex WUDataset. For the RoNIN-ds dataset, where motion patterns tend to be simpler and more repetitive, the accuracy improvements offered by ASLE are less dramatic. This is likely because existing baseline models already perform sufficiently well in mapping motion features to displacement within the context of the dataset. However, ASLE still achieves measurable gains, reinforcing its generalization capability. In general, these results highlight the effectiveness of ASLE in reducing displacement errors and its robustness to variations in input sequence length and complexity.
@@ V-C Experimental Results of ASLE Network
TABLE III: Computation overhead of ReNiL and other methods in RoNIN-ds test set.
Method Parms (M)
FLOPs (G) for
20s inference
MAE over
RoRes18
@@
ReNiL 16.41 3.33 91%
@@
TABLE IV: Comparison of device orientation estimation performance, the best result is marked with red.
Subset Metric ReNiL Mahony Madgwick
Vicon QAE 0.807 1.536 1.273
CS 0.733 0.350 0.538
Lidar-SLAM QAE 0.383 1.494 1.576
CS 0.837 0.429 0.329
Computation overhead
Table III summarizes the computation overhead for different methods in terms of parameter count, FLOPs, and localization accuracy (represented by MAE relative to RoRes18 in RoNIN-ds). PDR exhibits minimal computational complexity, but it yields significantly worst accuracy. RoRes18 and TLIO have similar parameter counts and computational overhead, while TLIO slightly improves accuracy to 98% MAE. IONet substantially reduces the size of the model and the computational cost, but achieves only 326% MAE. ReNiL, despite having the largest parameter count, demonstrates efficient inference along with the best localization accuracy. ReNiL improves accuracy while reducing FLOPs by leveraging the fast reasoning capabilities of convolutional networks and its ability to model any scale sequence. ReNiL has the potential to sacrifice some accuracy in exchange for a huge reduction in FLOPs, for example, by increasing the kernel size and stride of the TCN compressor in the embedding layer.
V-D Orientation Alignment Performance
Effectiveness performance
Fig. 10 shows the results of these methods and the ground truth for device orientation during one example. As shown in Table IV, our method significantly outperforms Mahony and Madgwick in estimating the orientation of the device during pedestrian movement, both when using Vicon and Lidar-SLAM systems. Our method converges the device orientation estimation framework to the navigation coordinate system in a short time, and maintains stability and effectiveness even under complex pedestrian motion patterns and magnetic field changes. In contrast, Mahony and Madgwick tend to experience orientation estimation fluctuations and heading drift due to the complexity of pedestrian motion patterns and magnetic field disturbances.
V-E Application Performance

### v3

@@ 4.1 Orientation Alignment
4.2 ASLE Structure
We propose an any-scale Laplace estimator (ASLE), which consists of five key components: 1) the patching module, 2) the input module, 3) the feature extractor, 4) the contextual builder, and 5) the regression module. Each module within the ASLE can be replaced with other general modules that perform the same function. In this paper, we have designed a specific verification model for the system. The structure of the ASLE is shown in Fig. 3.
@@ 4.2 ASLE Structure
Input module
The input module is used to map the original 6 features of the input $X\in\mathcal{R}^{B\times P\times 6\times L}$ to a higher-dimensional channels and compress the patch length to greatly reduce computation overhead. Specifically, a typical convolutional input layer is applied to embed $X$ from 6 channels into a higher channels and L into a shorter length:
@@
where $e$ is the channels dimension and $L^{\prime}$ is the compressed length.
@@
The 1D version of ResNet18 [he2016deep] has demonstrated strong sequence feature extraction capabilities in previous work. We use its improved residual block (replace BN with GN) as a feature extractor to capture motion pattern features in each patch in parallel. Specifically, for $X^{e}$, multiple 1D residual blocks are applied sequentially to extract patch features. Each residual block consists of two convolutional layers and two GN layers. The number of channels in each residual block is double that of the previous block. The module contains a total of $N_{1}$ residual blocks. For each block, the convolutional layer has a stride of 2 for first layer, 1 for others and a kernel size of 3. The output of each 1D residual block serves as the input for the next 1D residual block. The outputs of the final residual block are concatenated along the channel axis. The output of this module is:
@@
The multi-head self-attention is highly effective for context modeling and able to handle sequences of any scale. It also offers fast computation, making it an excellent choice for modeling movement context between patches. Specifically, to allow the attention layer to model features in the last two dimensions, we flatten the second and third axes of the output tensor $X$ from the 1D feature extractor, and add an additional zeros regression tokens to the time dimension. This yields a new tensor
$X\in\mathcal{R}^{B\times(P+1)\times(C\times W)}$ (16)
Then, $N_{2}$ multi-head self-attention blocks are applied to build the temporal local context of the features, resulting in a robust features map. Each multi-head self-attention block works as follows:
$\mathrm{MHA}(X)=\Big(\mathrm{softmax}\!\left(\frac{(XW_{Q})(XW_{K})^{\top}}{\sqrt{d}}\right)(XW_{V})+X\Big)W_{O}$ (17)
Where $W_{Q},W_{K},W_{V},W_{O}$ are learnable projection matrices, d is attention scaling dimension. Finally, we obtain robust features map:
$X\in\mathcal{R}^{B\times(P+1)\times D}$ (18)
Where $D$ is the attention dimension. This operation aims to fuse the contextual potential relationships between patches, extract the displacement information of the entire IMU into regression tokens, and improve the robustness of the final feature map.
Regression module
The regression module aims to regress the inherent features in regression tokens into Laplace parameter outputs. Specifically, a multilayer perceptron (MLP) $O(\cdot)$, which contains fully connected, dropout, and ReLU layers, is applied to regress $\Delta\hat{p},\hat{b}$.
@@
4.3 Joint Training with Uncertainty Estimation
@@
Specifically, for the original input $X\in R^{B\times P\times 6\times L}$ of the network, we record its features map $X_{f}\in\mathcal{R}^{B\times(P+1)\times D}$, which is the output after the contextual builder. We then perform a series of data augmentations on the original input $X$, including:
@@
Subsequently, we obtain the features map $\hat{X}_{f}\in\mathcal{R}^{B\times(P+1)\times D}$ of the contextual builder output. Finally, we calculate the MSE loss $L_{fm}$ between $\hat{X}_{f}$ and $X_{f}$ to construct robust context modeling features:
@@
4.4 Bayesian Process
The output of ASLE is the relative displacement and scale parameters between two IPDPs. However, our localization inference is a continuous process. Next, we will use Bayesian inference to connect all IPDPs in series to form a complete Bayesian inference chain. In the case of IPDPs involved with other external localization sources, we use a Bayesian filter [chen2003bayesian] for loose coupling.
@@ 4.4 Bayesian Process
$\mathbf{p}\left({p}_{t_{n}}\mid u_{t_{0}:t_{n}}\right)=\int\mathbf{p}\left({p}_{t_{n}}\mid{p}_{t_{n-1}},u_{t_{n}}\right)\mathbf{p}\left({p}_{t_{n-1}}\mid u_{t_{0}:t_{n-1}}\right)d{p}_{t_{n-1}}$ (27)
@@
$\mathbf{p}\left(\widetilde{p_{t_{n}}}\mid{z}_{t_{0}:t_{n-1}},u_{t_{0}:t_{n}}\right)=\int\mathbf{p}\left(p_{t_{n}}\mid p_{t_{n-1}},{u}_{t_{n}}\right)\,\mathbf{p}\left(p_{t_{n-1}}\mid{z}_{t_{0}:t_{n-1}},{u}_{t_{0}:t_{n-1}}\right)\,dp_{t_{n-1}}$ (29)
@@
5 Experimental Evaluations
5.1 Experiment Setup
@@
For displacement-regression experiments we employ ASLE-ns, a family of our ASLE that accepts fixed n-second input windows. Because the competing baselines PDR, RoNIN-ResNet18 (RoRes18), TLIO, CTIN, and iMoT can each operate only at a single preset window size, we instantiate four representative variants (n = 1, 5, 10, 20) to span short, medium, and longer motion segments; these values were chosen simply to give broad temporal coverage rather than for any dataset-specific reason. We evaluate all methods head-to-head and then carry forward the two strongest baselines for deeper analysis. For device orientation estimation, our motion-aware orientation filter is benchmarked against the classical Mahony [mahony2008nonlinear] and Madgwick [madgwick2011estimation] algorithms.
@@
The WUDataset collection was carried out using Google Pixel 6A, Huawei mate 40E, Iphone 16 Pro and Xiaomi 15 smartphone (IMU sampling rate: 200Hz), a Vicon optical system (sampling rate: 100Hz), and a LiDAR-SLAM system (sampling rate: 200Hz), all synchronized through a custom-developed Android application. The WUDataset is organized as follows:
@@
Table 1: Evaluation Metrics and Definitions
@@
MAE $\mathrm{MAE}=\frac{1}{N}\sum_{i=1}^{N}\big\|\hat{\bm{p}}_{i}-\bm{p}_{i}\big\|_{1}$
ADE $\mathrm{ADE}=\frac{1}{N}\sum_{i=1}^{N}\frac{\big\|\Delta\hat{\bm{p}}_{i}-\Delta\bm{p}_{i}\big\|_{2}}{\Delta t_{i}}$
HE $\mathrm{HE}=\frac{1}{N}\sum_{i=1}^{N}\big|\hat{\theta}_{i}-\theta_{i}\big|$
QAE $\mathrm{QAE}=\frac{1}{N}\sum_{i=1}^{N}2\,\arccos\!\big(\,|\langle\hat{\bm{q}}_{i},\bm{q}_{i}\rangle|\,\big)$
CS $\mathrm{CS}=\frac{1}{N}\sum_{i=1}^{N}\frac{\langle\hat{\bm{q}}_{i},\bm{q}_{i}\rangle}{\|\hat{\bm{q}}_{i}\|\,\|\bm{q}_{i}\|}$
5.2 Implementation Details
@@
ASLE details
We implemented ASLE using the PyTorch 2.5.1 [pytorch251] framework on a server equipped with an Intel Xeon Silver 4210R CPU and a GeForce GTX 3090 GPU. The model utilizes 200Hz IMU input, with a patch length of 200. The channel, kernel size, and stride of the input module are set to 32, 7, 2. The feature extractor consists of four 1D residual layers with output channels of 32, 64, 128, and 256, respectively. The context builder consists of 4 MHA blocks with 4 heads, and 64 dimensions for each patch. The regression module is conducted as two fully connected layers with hidden size 128, ReLU layer, and the output is a vector of length 4. Adam [kingma2014adam] is used as the optimizer, with the learning rate scheduler set to ReduceLROnPlateau [pytorch_reduce_lr_on_plateau], using a factor of 0.1 and patience of 10. The initial learning rate for the network parameters is set to $1\times 10^{-4}$, with a minimum learning rate of $1\times 10^{-8}$. The mini-batch size is 32, and the model is trained for a total of 200 epochs.
Bayesian process
$\tau_{t_{n}}\sim\operatorname{Exp}(\frac{1}{2\bm{\Sigma}_{b}^{2}})$ (31)
where $\bm{\Sigma}_{b}=\operatorname{diag}\left(\Delta t_{n}b_{v_{x}},\Delta t_{n}b_{v_{y}}\right)$. So the Laplace distribution noise can be conditionally Gaussian representation into a Gaussian distribution noise:
${w}_{t_{n}}\mid\tau_{t_{n}}\sim\mathcal{N}\left({0},\tau_{t_{n}}\bm{\Sigma}_{w}\right)$ (32)
where $\bm{\Sigma}_{w}=\operatorname{diag}\left(2\left(\Delta t_{n}b_{v_{x}}\right)^{2},2\left(\Delta t_{n}b_{v_{y}}\right)^{2}\right)$. Then, we use $\tau_{t_{n}}\bm{\Sigma}_{w}$ to construct a covariance matrix that conforms to the Gaussian distribution to complete the Bayesian inference chain.
$\delta_{t_{n}}=\left(p_{t_{n-1}}-p_{t_{n}}\right)^{\top}\bm{\Sigma}_{w}^{-1}\left(p_{t_{n-1}}-p_{t_{n}}\right)$ (33)
5.3 Experimental Results of ReNiL Framework
Displacement accuracy performance
Table 2: Displacement accuracy on RoNIN-ds and WUDataset. Top-1 bold red, Top-2 bold blue, Top-3 bold green. Ties share the same rank.
Dataset Subjects Metric $\downarrow$ PDR RoRes18 TLIO CTIN iMoT ASLE-1s ASLE-5s ASLE-10s ASLE-20s
@@ 5.3 Experimental Results of ReNiL Framework
The purpose of this experiment is to evaluate the displacement accuracy of the ASLE network compared to several baseline methods, tested on both seen and unseen subjects in the RoNIN-ds and WUDataset datasets. As shown in Table 2, when limited to a fixed time scale of 1 second, the displacement accuracy of all methods remains relatively similar across both datasets. When the time scale restriction is removed, the distinctive advantage of ASLE becomes evident. Its inherent design to process variable-scale IMU sequences allows it to outperform other methods that rely on fixed-length sliding windows. This flexibility enables ASLE to effectively capture temporal dependencies, which translates into substantially improved displacement accuracy, particularly pronounced on the more complex WUDataset. For the RoNIN-ds dataset, where motion patterns tend to be simpler and more repetitive, the accuracy improvements offered by ASLE are less dramatic. This is likely because existing baseline models already perform sufficiently well in mapping motion features to displacement within the context of the dataset. However, ASLE still achieves measurable gains, reinforcing its generalization capability. In general, these results highlight the effectiveness of ASLE in reducing displacement errors and its robustness to variations in input sequence length and complexity.
Alleviation of integration error
@@
Cross collector, device, and pattern applicability performance
Table 3: Cross-split evaluation under different grouping strategies in WUDataset (10s window).
Grouping strategy Training split Test split Mean MAE Variance
Train Test Train Test
Collectors Collectors G1 Collectors G2 1.35 1.74 0.002 0.002
Patterns Patterns G1 Patterns G2 1.29 1.91 0.002 0.004
Phones Phones G1 Phones G2 1.41 1.69 0.001 0.002
Collectors and patterns Collectors G1 + patterns G1 Collectors G2 + patterns G2 1.26 1.97 0.003 0.004
Phones and patterns Phones G1 + patterns G1 Phones G2 + patterns G2 1.31 1.89 0.002 0.003
Collectors and phones Collectors G1 + phones G1 Collectors G2 + phones G2 1.34 1.71 0.001 0.002
In this section, we employ multiple grouping strategies to conduct experiments. G1 and G2 involve randomly dividing 28 collectors, 9 patterns, and 4 device types into two uniform, disjoint parts according to the grouping strategy. Collectors G1 G2 means dividing the dataset into two disjoint, uniform parts based on collectors, and then using one part for training and the other for testing. Collectors G1 + patterns G1, Collectors G2 + patterns G2 means using the data collected by Collectors G1 with patterns G1 for training, and using the data collected by Collectors G2 with patterns G2 for testing.
In Table 3. ReNiL demonstrates strong applicability across different collectors, devices, and patterns. It maintains consistent performance, with only moderate variations in mean absolute error (MAE) and variance, even when trained on one set of conditions and tested on another. Specifically, the model shows minimal error across different collectors (1.35 vs. 1.74), devices (1.41 vs. 1.69), and patterns (1.29 vs. 1.91), highlighting its robust cross-collector and cross-device adaptability. Additionally, when combining multiple factors, such as collector and pattern combinations, or device and pattern combinations, the model continues to deliver stable results, with only slight increases in error and variance. This indicates the model’s high versatility and robustness in real-world, multi-user, multi-device, and diverse pattern scenarios.
Computation overhead
Table 4: Computation overhead of ReNiL and other methods in RoNIN-ds test set.
Method Parms (M) FLOPs (G) for 20s inference MAE over RoRes18
@@
ReNiL 2.09 0.38 91%
Table 4 summarizes the computation overhead for different methods in terms of parameter count, FLOPs, and localization accuracy (represented by MAE relative to RoRes18 in RoNIN-ds). PDR exhibits minimal computational complexity, but it yields significantly worst accuracy. RoRes18 and TLIO have similar parameter counts and computational overhead, while TLIO slightly improves accuracy to 98% MAE. IONet substantially reduces the size of the model and the computational cost, but achieves only 326% MAE. ReNiL has 33.3% the number of parameters and about 4.6% the FLOPS of RoRes18, and exhibits efficient reasoning capabilities and optimal positioning accuracy. ReNiL improves accuracy while reducing FLOPs by leveraging the fast reasoning capabilities of convolutional networks and MHA’s ability to model any scale sequence.
Resource usage comparison
Figure 10: An example of a 100-second positioning process running on a mobile phone with 10 positioning requests. Comparison of resource usage between ReNiL’s ASLE and fixed-rate 5HZ RoRes18.
Fig.10 shows an example of a 100-second positioning process running on a mobile phone, consisting of 10 positioning requests. ReNiL significantly reduces memory usage, utilizing device memory and computing power almost exclusively at the IPDPs. In contrast, conventional trajectory tracking algorithms maintain high levels of sustained memory usage while linearly increasing FLOPs. Clearly, even in the worst-case scenario, where the application requirement is simply tracking user trajectory, ReNiL still outperforms with a smaller model size and lower single-inference cost.
Orientation effectiveness performance
Table 5: Comparison of device orientation estimation performance, the best result is marked with bold red.
Subset Metric ReNiL Mahony Madgwick
Vicon QAE $\downarrow$ 0.807 1.536 1.273
CS $\uparrow$ 0.733 0.350 0.538
Lidar-SLAM QAE $\downarrow$ 0.383 1.494 1.576
CS $\uparrow$ 0.837 0.429 0.329
@@
Fig. 11 shows the results of these methods and the ground truth for device orientation during one example. As shown in Table 5, our method significantly outperforms Mahony and Madgwick in estimating the orientation of the device during pedestrian movement, both when using Vicon and Lidar-SLAM systems. Our method converges the device orientation estimation framework to the navigation coordinate system in a short time, and maintains stability and effectiveness even under complex pedestrian motion patterns and magnetic field changes. In contrast, Mahony and Madgwick tend to experience orientation estimation fluctuations and heading drift due to the complexity of pedestrian motion patterns and magnetic field disturbances.
5.4 Application Performance in Real World